OpenStudy (anonymous):

Determine the range of the function f(x)=3x-9 when the domain is {-2, 0, 3}. {-6, 0, 9} {-15, -9, 0} {3, 9, 18} {-2, 0, 3}

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

The domain is the set of x values for the function f(x)=3x-9. So, you have to crank out the y values for the function using the given x values. To help you get started, I'll do the first one. f(x)=3x-9 f(-2) = 3 * (-2) - 9 = -6 -9 = -15.

Directrix (directrix):

-15 is one of the three values of the range. If x = 0, then f(0) = what? @Demi_xx

OpenStudy (anonymous):

my answer is b yea?

Directrix (directrix):

{-15, -9, 0} is corrrect.